OpenStudy (anonymous):

2. If a stone is tossed from the top of a 250 meter building, the height of the stone as a function of time is given by h(t) = -9.8t2 – 10t + 250, where t is in seconds, and height is in meters. After how many seconds will the stone hit the ground? Round to the nearest hundredth’s place; include units in your answer.

OpenStudy (anonymous):

Calc course?

OpenStudy (anonymous):

The type of course this is affects the method for finding the answer.

OpenStudy (anonymous):

what you mean

OpenStudy (zzr0ck3r):

-9.8t^2-10t+250 = 0

OpenStudy (zzr0ck3r):

do you know what to do from there?

OpenStudy (anonymous):

Ah, nevermind. Didn't read the question carefully. No calculus necessary.

OpenStudy (anonymous):

zzr0cker has it right. When it hits the ground, its height is 0. So set the height equal to 0 and solve for t.

OpenStudy (anonymous):

make h=0 because h at ground is 0 and calculate time neglect -ve time from the 2 roots

OpenStudy (anonymous):

what

OpenStudy (anonymous):

4.57 sec

OpenStudy (anonymous):

but how

OpenStudy (anonymous):

because i want to learn it

OpenStudy (anonymous):

Okay, so they give us an equation for height based on time, right? Do you agree that the height when it hits the ground is 0?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Cool, so let's set our equation for height to 0 and figure out what value of t makes it true. Like this: -9.8t^2 - 10t +250 = 0 solve for t. And do you remember how to solve an equation of the form Ax^2 + Bx + C = 0?

OpenStudy (anonymous):

no i drop out of school just recenlty enroll back in sorry

OpenStudy (zzr0ck3r):

there are many ways, what are you guys doing in class? Does A C method sound familiar? or completing the square? or (-b +- sqrt(b^2-4ac))/2a

OpenStudy (anonymous):

It's okay. There's this cool formula that looks complicated and fancy, but is actually pretty easy to use. It's called the quadratic formula and it works whenever we have an equation like Ax^2 + Bx + C = 0. Figure out what A, B, and C are and plug them in to solve for x: \[\frac{-B \pm \sqrt{B^2-4AC}}{2A}\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

What are A, B, and C?

OpenStudy (zzr0ck3r):

you could also graph it and see where it crosses the x-axes it all depends on what your teacher wants

OpenStudy (anonymous):

wait am still lost

OpenStudy (phi):

Here is a video, if you have time http://www.khanacademy.org/math/algebra/quadtratics/v/introduction-to-the-quadratic-equation

OpenStudy (anonymous):

Let me show you a simple example of myself using the formula. Suppose I'm trying to solve the equation \[2x^2 +2x -12 = 0\] First, I notice that the equation looks like \[Ax^2 + Bx + C =0\] With A = 2, B=2, and C =-12 So, that means to solve, I can plug those into: \[\frac{-B \pm \sqrt{B^2-4AC}}{2A}\] So, I get: \[\frac{-2 \pm \sqrt{2^2 -4*2*(-12)}}{2*2}\]

OpenStudy (anonymous):

Doing a little bit of math on that gives me: \[\frac{-2 \pm \sqrt{4-(-96)}}{4}\] which gives \[\frac{-2 \pm \sqrt{100}}{4}\] so I get \[\frac{-2 \pm 10}{4}\]

OpenStudy (anonymous):

I get two answers from that. One comes from using a plus sign \[\frac{-2+10}{4} = \frac{8}{4} = 2\] The other comes from using a minus sign \[\frac{-2-10}{4} = \frac{-12}{4} = -3\]

OpenStudy (anonymous):

So that tells me, with just a little bit of work, that the solutions to my original equation are x=2, and x=-3. The cool thing about the equation is that it ALWAYS works! =D

OpenStudy (anonymous):

So, for your problem -9.8t^2 -10t +250 = 0 it's the same exact process, but A, B, and C are different numbers, and your answers won't be such pretty integers like I got. They'll be decimals.

OpenStudy (anonymous):

Oh, and you'll get one positive answer and one negative answer. The negative answer says that it hits the ground BEFORE he dropped it, which doesn't make any sense. So we ignore that one. The positive answer is correct though. Give it a try, and ask me any questions you have.

OpenStudy (anonymous):

Are you allowed to use calculus? It would ba a lot easier

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Actually, Naomi, since it doesn't ask for the maximum height of the projectile, but just when it lands, it really doesn't require any calculus.

OpenStudy (anonymous):

o

OpenStudy (anonymous):

I'm merely suggesting another way he could do it.

OpenStudy (anonymous):

i dont understand

OpenStudy (anonymous):

Do you know how to get derivatives? If not then follow Smoothmath's method without calculus

OpenStudy (zzr0ck3r):

I dont see how setting a function equal to zero can get any easier, how would you do this in a "easier" way with calc?

OpenStudy (anonymous):

I'm with zz. I don't think derivatives actually do anything for you in this problem.